verapamil HCI 0.075 mg/kilogram IV push over two minutes with the label on the pile reads 5 mg/2 ML and the patient weighs 40 kg how many Milla liters/MIN will you administer
Answer #1

ANSWER.

* Calculation of amount(ml) of drug required.

         Formula = Desired dose/Dose in hand * Quantity.

           Ordered dose = 0.075mg/kg.

           Weight of the patient = 40 kg.

         So

            Desired dose(D) = 0.075 * 40

                                        = 3 mg.

            Dose in hand (H) = 5 mg.

                       Quantity(Q) = 2 ml.

            As per formula

                 Amount(ml) of drug required = D/H *Q

                                                               = 3/5 * 2

                                                               = 0.6 * 2

                                                               = 1.2 ml per 2 minutes.

       Amount of drug administered per minute = 1.2/2

                                                                         = 0.6 ml.